OpenStudy (kise_ryouta):

Solve for a. 4(a + 3)=12 + 4a4(a + 3)=12 + 4a no solution all real numbers −3−3 −1

OpenStudy (benji):

Alright before we can bring all the a's to one side, you have to distribute the brackets, what is \[4(a+3)\]distributed?

OpenStudy (kise_ryouta):

4a+12

OpenStudy (danieldjpon3):

Yes

OpenStudy (danieldjpon3):

Correct!

OpenStudy (benji):

Exactly so now we have \[4a+12=12+4a\]What do you think is up next?

OpenStudy (kise_ryouta):

comb like terms?

OpenStudy (kise_ryouta):

*combine

OpenStudy (benji):

Yep, something weird happens when you try to do that though, right?

OpenStudy (benji):

Let's combine like terms, how do you bring the 4a to one side?

OpenStudy (kise_ryouta):

subtract right

OpenStudy (benji):

Yes, so we have \[4a-4a+12=12+4a-4a\], and the 4a's cancel! So do the 12's! there is no solution.

OpenStudy (kise_ryouta):

THANKS KOOKIE BTS FOR LIFE

OpenStudy (benji):

lol

OpenStudy (kise_ryouta):

can you help me with another one